LUCIFER Com2
UT 2008 December 05

Observers: N.Ageorges, W.Seifert, P.Buschkamp, M.Juette, V.Naranjo, several others from LUCIFER team. D.Thompson (@LBT)
Telescope Operator: D.Gonzalez-Huerta
SW Support: N.Cushing, T.Leibold in Tucson until...

Data will eventually be copied to /Repository/LUCIFER_Data/20081204/

Summary:

  • Started cloudy, but it cleared off. Opened at 09:16 and worked the remainder of the night.
  • GCS crashes cause by acquisition algorithm, fixed but patch to wait until tomorrow
  • Huge IE/CA offsets from yesterday midnight, encoders needed resetting
  • Rotator offsets work in RADEC space but not in DETXY?
  • Several "Chase failed to make tolerance"s recorded by LUCIFER as errors but system continued working.
  • Stable, good seeing, but again failed to obtain focus measurement for incorrect setfocus command (even though it claimed to work) and preference to complete some other LUCIFER tests tonight.
  • Troubles at end of night caused restarts of GCS, GCSGUI, IIF, IRS, and a client(?) on LUCIFER side.
  • Many jumps seen in LUCIFER data (see below).

Details:

09:16 Dome open! H=70% and has been climbing over the last several hours but seems to have plateaued.

09:34 First active preset failed with a crash of GCS. Acquire000054.fits does show a reasonably centered source with low peak counts (~40DN over background).

09:54 Reloaded the default pointing model and the pointing star was found ~45 arcsec offset from center, mostly in IE. Last night we were ~100 arcsec offset in the other direction. Outside temp is -2.9, several degrees cooler than last night.

10:00 Preset after correcting pointing failed and pointing was off by 16 arcsec.

10:49 Again large offset to pointing. Odd that it is mostly in CA. Currently IE=-83.1 CA=19.35, these are 160 arcsec in IE and 80 in CA off from ~midnight last night. Outside temps still flat at -2.9C. We now suspect that the encoders need to be reset, but the LUCIFER team does not want to take the time right now to run the encoder reset.

01:42 LUCIFER team conducting offset tests for the last ~3 hours, same field.

01:42 Started encoder reset run. Re-loaded 20081017 pointing model. Done 01:54

02:00 PTModel adjusted to IE=-14.91 CA=16.35 (offset only in CA, ~26 arcsec). LUCIFER continuing with their offset tests.

02:21:32.278 "MirrorGoToPositionAbs Command Failed because "Chase failed to make tolerance" is reported as an error to LUCIFER but guiding and WFSing continue. Issue Trak 1451, 1891 report essentially that this is an occasional problem, more pronounced in colder weather. Added note to IT#1891.

We are getting many jumps (see attached image). This is six coadded 5s integrations immediately after an offset. The subsequent 30s summed image also shows a single jump. We are running in "default" mode with coma and focus going to the secondary.

  • Jumps seen in LUCIFER images:
    jumps.jpg

Times from JMH relevant to the above event

stepFocus finished at 02:47:45
guiding at 02:47:51
guiding at 02:47:57
guiding at 02:48:03
guiding at 02:48:09
guiding at 02:48:16 - star well centered in guider 6285
wfs correction at 02:48:26
guiding at 02:48:27 - 1 arcsec correction from guider 6286 (4.6 sec)
M1 starts to move 02:48:28
M1 finishes moving 02:48:33
guiding at 02:48:33 - another 1 arcsec correction from guider 6287 - this one could have been influenced by M1, but star hasn't moved since 6286
M2 moves 02:48:34
guiding continues until.....
guiding at 02:449:22

next offset at 02:49:23

(Something made the image move by an arcsec, perhaps 2 arcsec, but I can't explain what it was. - JMH)

End of JMH notes

02:53:55 Same "Chase" message again

03:02 Taking some pupil images.

03:19 Resuming offset tests. Note: rotator offsets in offset pointing requests work in OFFSET_RADEC space but not in OFFSET_DETXY space (the XY offsets to work as expected in both systems). Email sent to Michele to clarify. See also IT#1611.

Additional "Chase" errors at 03:53 and 03:58

04:22 Seeing is good and seems stable, try another AGW-LUCIFER focus sequence.
AIP_L.cfg shows focus=31.6mm, guider is showing 0.45 arcsec FWHM.

  • Data: luci_20081204_
    LUCI# Focus FWHM
    0132 25.359 3.14
    0133 19.359 way out of focus
    0134 22.359  

Aborted sequence. The focus stage for some reason went to 0.0mm even though the request to move to 25.359 claimed to have completed successfully. Re-sending the preset to re-establish collimation, then starting focus sequence over. The same behavior repeated a second time. We re-homed the stages and then simply re-sent the preset. Oops...my fault. The problem originated because I sent the setfocus without a -p, just the number.

04:56 Third try worked, of course..but...

  • Data: luci_20081204_
    LUCI# Focus FWHM last WFS rms
    0136 25.355 4.53 156.72
    0137 19.356 3.39 81.76
  • Results: WFS would not restart from the GCSGUI either on lbtmu05 or the TO station. Aborting the focus sequence because of other pending tests the LUCIFER team wants to finish before dawn.

05:20 Continuing with other LUCIFER tests. GCSGUI is still acting oddly, but GCS is guiding and WFSing. The guider is seeing ~0.4 arcsec FWHM and the WFS is staying well below 150nm rms.

05:27:23 A setfocus 0.0 sent from LUCIFER hung and did not time out after >5 minutes. IIF was killed and restarted, along with GCS and GCSGUI because they were having trouble

12:40:51 Preset failed for lack of guide star on acquire00069.fits. There is a guide star at 8000 DN above the background and well focused. It is very close to the bias region. Norm thinks that the patch that Torsten already prepared would fix this.

05:42:47 Preset sent in TRACK mode, succeeded. 05:43:19 StepFocus sent, failed with "invalid buffer"

06:46 End of night: Closing

-- DavidThompson - 05 Dec 2008
Topic attachments
I Attachment Action Size Date Who Comment
jumps.jpgjpg jumps.jpg manage 10 K 05 Dec 2008 - 10:08 DavidThompson Jumps seen in LUCIFER images
Topic revision: r3 - 06 Dec 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback